OpenStudy (anonymous):

Metal will be awarded!!!! In what quadrant does the solution of the equation fall? y = x - 1 y = -3x - 5 Answers 1. 2. 3. 4.

OpenStudy (welshfella):

ebcuase both right sides are equal to x we can write x - 1 = -3x - 5 add 3x to both sides 4x - 1 = -5 can you continue to find the value of x?

OpenStudy (welshfella):

* are equal yo y

OpenStudy (anonymous):

-1 = 9?

OpenStudy (anonymous):

or -1 = 1?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

you have to have x in the equation x = -1 4x - 1 = -5 add 1 to both sides 4x -1 + 1 = -5 + 1 4x = -4 divide both aides by 4 4x -4 -- = -- 4 4 x = -1

OpenStudy (welshfella):

now to find the value of y plug x = -1 into y = x - 1:- y = -1-1 = ?

OpenStudy (welshfella):

what is -1-1 equal to?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

No -1-1 = -2 

OpenStudy (anonymous):

so what quadrant

OpenStudy (welshfella):

so the solution of system of equations is x = 1 , y = -2 that is the point (-1,-2)

OpenStudy (welshfella):

so which quadrant is that in?

OpenStudy (anonymous):

so the second quadrant?

OpenStudy (welshfella):